# LiSA: Leveraging Link Recommender to Attack Graph Neural Networks via Subgraph Injection

### v1

@@ 4.2. Training Algorithm
The overall training algorithm of LiSA is shown in Algorithm 1. The training procedure for the subgraph involves a two-tiered process with inner and outer iterations. In the inner iterations, we simultaneously train surrogate models for link prediction and node classification to optimize model parameters, while the outer iterations optimize the subgraph by alternately updating its adjacency matrix and node features. The inherent complexity of graph-structured data means that changes to the edges can significantly impact node representation. This alternating update methodology is crucial, as simultaneous modifications to both the adjacency matrix and node features can lead to instability and challenges in achieving convergence.
@@
4.2.1. Feature Optimization
The optimization process for refining the features within the subgraph employs a conventional approach that utilizes the principles of gradient descent through backpropagation. The goal is to adjust the node features in a manner that directly contributes to the two adversarial objectives. Each iteration of feature optimization is guided by the following update rule:
(14) $F_{s}=F_{s}-lr\cdot\nabla_{F_{s}}\mathcal{L}_{atk}$
4.2.2. Structure Optimization
(15) $\nabla_{str}=\nabla_{A_{p}}\mathcal{L}_{atk}^{link}+\beta\cdot\nabla_{\hat{A}}\mathcal{L}_{atk}^{cls}$
In this formula, $\nabla_{A_{p}}\mathcal{L}_{atk}^{link}$ denotes the gradient of attack loss by the link prediction model with respect to the subgraph adjacency matrix $A_{p}$ before link recommender intervention, while $\nabla_{\hat{A}}\mathcal{L}_{atk}^{cls}$ represents the gradient of attack loss by the node classification model with respect to the post-update adjacency matrix $\hat{A}$ modified by the link recommender. The parameter $\beta$ acts as a balancing factor in the optimization of the structure.
The structural update process involves ranking the combined gradient $\nabla_{str}$, then strategically adding and removing edges based on this ranking. Specifically, we add edges in locations where they are currently absent, prioritizing those with the lowest combined gradient values. Simultaneously, we remove edges from areas with the highest combined gradient values:
(16) $A_{s}=A_{s}+\delta(A_{s},\min(\nabla_{str}))-\delta(A_{s},\max(\nabla_{str}))$
Here, $\delta()$ is the function to add or remove edges from an adjacency matrix according to the combined gradient ranking. This iterative modification ensures the adjacency matrix is finely tuned to meet the adversarial objectives, enhancing the capability of the subgraph to disrupt both link prediction and node classification models.
5. Experiments
In this section, we evaluate our methods on various datasets to answer the following research questions:
  • RQ1: Can our method deceive a link recommender into generating links between target victim nodes and the subgraph while also compromising the classification accuracy of target nodes?
  • RQ2: How does the subgraph size affect the effectiveness of LiSA? What impact does the quantity of links added by users have on the attack performance?
  • RQ3: To what extent is each building block of LiSA indispensable for facilitating a successful attack? How does each component enhance the overall effectiveness of LiSA?
5.1. Experiment Settings
5.1.1. Datasets
To evaluate the effectiveness of our attack methods, we perform experiments on four distinct datasets, each offering a unique combination of graph size and average node degree to test our methods across different scenarios. These datasets include Cora and PubMed, both citation networks, Amazon-Photo and Computers, co-purchase networks from a popular e-commerce platform, and FacebookPagePage, representing a social network graph. Detailed statistics of these datasets are summarized in Appendix B.
5.1.2. Experimental Verification
In this study, we conduct an experimental verification of our LiSA framework using several prominent GNN models. These models include GAE and VGAE for link recommender, along with GCN, SGC, and GraphSAGE as node classifiers. Our attack methodology is applied across all datasets with the presumption that the target node would establish links with nodes of top three linking scores in the entire poisoned graph. To assess the feasibility of our methods, we randomly select 1000 nodes from each original graph as target victim nodes. We then generate subgraphs and execute attacks on these nodes individually. This approach allows us to compute both the link success rate (LSR) and the overall attack success rate (ASR). The LSR is defined as the proportion where at least one adversarial edge is generated between the target node and subgraph, while the ASR refers to the misclassfication rate on these victim nodes. By examining the efficacy on each target node separately, we avoid interference among the attached nodes, providing a fair and isolated assessment of each attack. For the link recommender, we use 85% of the edges for training, with half serving as positive edge labels with negative sampling for both subgraph optimization and attack verification. For node classification, 5% of the nodes are randomly chosen for training labels. These labels are utilized in both subgraph optimization and attack verification, accompanied by the pretraining of a node classifier to identify the second highest output class as the attack label. To maintain a consistent basis for comparison across different datasets, we employ a subgraph comprising five nodes for each target victim node attack. However, we tailor hyperparameters such as the number of subgraph edges $n_{E}$, and the balance parameters $\alpha$ and $\beta$ for each dataset, ensuring optimal performance and adaptability to the specific characteristics of each dataset (The detailed parameter settings can be found in the script). We also compare the LiSA framework with GraphCopy in the SIA scenario, and an additional baseline of conventional NIA scenario. This particular version of NIA incorporates the probability of successful link creation, aiming to demonstrate the influence of link generation on the success of NIAs. For additional experimental details and outcomes, please refer to Appendices A and C, respectively.
Figure 3. Attack outcome on GCN with SIA and NIA scenario. For SIA, the link recommender model is GAE.
5.2. Attack Results
In this subsection, we present the results of our attack methods to address RQ1. The attack results using the LiSA framework are presented in Table 1, while the misclassification rates for the original clean graphs are also provided, serving as a reference for evaluating the impact of our attacks. The outcomes of the attack indicate that the subgraphs produced by the LiSA framework successfully establish links between target nodes and subgraphs, leading to a notable decrease in the accuracy of node classification performance. This demonstrates the effectiveness of the LiSA framework within the SIA scenario. As illustrated in Fig. 3, the LiSA framework surpasses GraphCopy by achieving a higher ASR, thereby demonstrating its superiority over GraphCopy as an optimization-based method. Notably, the average attack budget for GraphCopy is proportional to the square of the average node degree in the dataset. In denser graphs, such as Amazon-Photo and Computers, attacking a single node requires a subgraph size exceeding 10% of the original graph. Conversely, LiSA achieves higher ASR with a much smaller budget, requiring only five nodes for each target node attack, highlighting the efficiency and practicality in real-world scenarios.
@@ 5.2. Attack Results
Table 1. LiSA attack results (LSR % — ASR %) and clean graph misclassification rate (%).
  Link Recommender Node Classifier Dataset
Cora Amazon-Photo Amazon-Computers PubMed FacebookPagePage
  - GCN 18.7 7.7 14.7 21.6 11.7
@@
  GAE GCN 52.3 — 48.6 31.0 — 30.4 34.0 — 35.5 82.3 — 68.5 50.4 — 48.0
GAE SGC 52.1 — 56.1 34.7 — 55.5 37.7 — 43.5 79.9 — 68.8 51.6 — 61.6
GAE GraphSAGE 52.6 — 55.1 34.5 — 29.2 35.8 — 32.6 80.2 — 64.3 49.6 — 47.0
VGAE GCN 44.6 — 50.3 25.2 — 29.5 32.9 — 37.2 77.3 — 66.8 42.4 — 45.0
VGAE SGC 45.9 — 59.9 26.2 — 58.2 33.9 — 43.1 79.5 — 68.2 35.8 — 52.6
VGAE GraphSAGE 46.8 — 55.5 26.8 — 21.8 35.6 — 32.5 79.4 — 67.4 40.2 — 37.4
5.3. Impacts of Subgraph Size and User Added Links
To answer RQ2, we conduct a series of experiments on the Amazon-Photo dataset. These experiments aim to investigate how changes in subgraph size and the number of links, chosen based on top linking scores, affect the LSR and ASR. Our initial observation reveal a decline in LSR as the subgraph size increased, a trend depicted in Fig. 4(a). This decline likely stems from the challenges associated with optimizing larger subgraphs, especially when adjusting both structure and features. However, we noted a stabilization in LSR when the subgraph size exceeds 5. Another interesting finding related to the response of ASR to the number of links added by users, as illustrated in Fig. 4(b). With the addition of a single link, we observe that the ASR is highest when the subgraph consist of a single node, decreasing as the subgraph size expands, possibly due to the corresponding reduction in LSR. On the other hand, with five links added, the ASR improve as the subgraph grow to five nodes, before experiencing a slight decrease with further increases in subgraph size. This phenomenon can be attributed to the rise in aggressor nodes linked to the target, enhancing effectiveness of the attack. However, exceeding the subgraph size with available links prompts the link recommender to seek top-scoring nodes from the original graph for link formation, which strengthens the resilience of target node to the attack (Dai et al., 2022a). Taking into account these effects, we found that the optimal ASR of 10 links was attained when the subgraph size was 7.
(a) LSR (%)
(b) ASR (%)
Figure 4. Impacts of subgraph size and user added links on Amazon-Photo.
5.4. Ablation Studies
To answer RQ3, we delve into the distinct contributions of each component of LiSA on the Cora and PubMed dataset, examining the effects on attack performance when specific elements are excluded, as depicted in Fig. 5. This analysis involves experiments with variations of the LiSA framework, including the omission of the surrogate models for node classification and link prediction (denoted as w/o cls and w/o link, respectively). Additionally, we explore the impact of focusing solely on node features optimization (labeled as w/o str) and solely on adjacency matrix optimization with randomly initialized node features (labeled as w/o feat) to discern the specific contributions of these optimization strategies.
(a) Cora
(b) PubMed
Figure 5. Ablation studies.
These findings reveal that adversarial links can be formed effectively through the use of the link prediction surrogate model, with the subgraph thus generated showing a marginal ASR improvement against the clean graph. Conversely, relying only on the node classification surrogate model fails to generate any links between target node and the subgraph, rendering it ineffective in contributing to the attack outcome. This highlights the significance of utilizing dual surrogate models, which can greatly enhance the attack to accomplish both adversarial objectives. Regarding optimization strategies, we observe that focusing exclusively on the structural aspects without adjusting the node features has no improvement on the attack. However, optimizing only the node features partially achieve the adversarial objectives, indicating that feature optimization plays a primary role in the SIA scenario. Nevertheless, structural optimization emerges as a critical supporting factor, enhancing the overall efficacy of the attack. Thus, the experimental results suggest that each component of LiSA is crucial to improve the attack performance.
5.5. Subgraph Analysis
(a) Cora
(b) PubMed
Figure 6. Subgraph Analysis.
5.6. Parameter Sensitivity Analysis
(a) LSR (%)
(b) ASR (%)
Figure 7. Hyperparameter sensitivity analysis.
5.7. Subgraph Training Time Analysis
Table 2. Subgraph training duration.
  Dataset Nodes Edges Epochs Duration (sec)
  Cora 2,708 10,556 200 7.47
Photo 7,650 238,162 300 215.47
Computers 13,752 491,722 300 650.18
PubMed 19,717 88,648 200 625.38
PagePage 22,470 342,004 200 1,036.42
Table 2 presents the subgraph training time of LiSA for each attack. The subgraphs of Cora/PubMed/FacebookPagePage, and Amazon-Photo/Computers were trained with 200 and 300 epochs, respectively, using an A6000 GPU with 48GB of memory. For LiSA, we utilized the adjacency matrix for optimization, resulting in a quadratic increase in training time with graph size. This is the primary limitation of the current version of our framework. We will explore an edge-weight-based approach and verify our method on large scale graphs as part of future work.
6. Conclusions
This study presents a new attack scenario, along with a novel attack framework that utilizes link recommender to compromise the node classification accuracy of GNNs through subgraph injection. Unlike conventional attacks that require direct manipulation of the original graph or the addition of links to artificially created nodes, LiSA executes an attack indirectly by injecting an isolated subgraph, thereby enhancing practicality in real-world scenarios. Our experiments across various real-world datasets demonstrate the effectiveness of LiSA, illustrating significant degradation in GNN performance. This work not only highlights a critical vulnerability in GNNs coupled with link recommenders but also sets a foundation for developing more robust GNN architectures.
Acknowledgements.
Appendix A GraphCopy Settings and Attack Results
GraphCopy introduces a fundamental approach for conducting adversarial attacks, achieving the adversarial objectives to a certain degree, as demonstrated in Table 3.
In the experimental configuration of GraphCopy, we create the subgraph by replicating the 2-hop neighborhoods surrounding the target nodes and introducing a 10% Gaussian perturbation to the features of the duplicated nodes. Detailed parameter settings are available in the script.
Table 3. GraphCopy attack results (LSR % — ASR %) and clean graph misclassification rate (%).
  Link Recommender Node Classifier Dataset
Cora Amazon-Photo Amazon-Computers PubMed FacebookPagePage
  - GCN 18.7 7.7 14.7 21.6 11.7
- SGC 18.7 50.1 44.0 25.3 26.8
- GraphSAGE 21.4 7.4 13.8 22.5 15.1
  GAE GCN 13.3 — 25.8 30.5 — 17.7 45.3 — 29.1 61.8 — 44.2 27.4 — 30.8
GAE SGC 14.0 — 35.5 30.8 — 56.0 44.4 — 46.3 62.6 — 51.7 29.0 — 56.0
GAE GraphSAGE 15.3 — 33.0 31.5 — 12.9 43.9 — 17.8 63.3 — 41.1 27.6 — 20.6
VGAE GCN 29.9 — 36.8 40.7 — 17.6 48.8 — 29.8 64.1 — 44.7 50.0 — 25.4
VGAE SGC 30.3 — 42.6 39.4 — 56.8 50.9 — 45.9 65.1 — 49.7 44.8 — 50.2
VGAE GraphSAGE 31.1 — 39.9 39.8 — 14.1 50.8 — 20.5 63.6 — 42.7 43.6 — 23.0
Appendix B Dataset Statistics
@@
Table 4. Dataset statistics.
@@
Amazon-Photo 7,650 23,8162 31.13 746 8
Amazon-Computers 13,752 491,722 35.76 767 10
@@
FacebookPagePage 22,470 342,004 15.22 128 4
Appendix C NIA Settings, Results and Analysis

### v2

@@ 4.2 Training Algorithm
The training procedure for the subgraph involves a two-tiered process with inner and outer iterations. In the inner iterations, we simultaneously train surrogate models for link prediction and node classification to optimize model parameters, while the outer iterations optimize the subgraph by alternately updating its adjacency matrix and node features. The inherent complexity of graph-structured data means that changes to the edges can significantly impact node representation. This alternating update methodology is necessary because it helps the training to achieve convergence.
4.2.1 Feature Optimization
The optimization process for refining the features within the subgraph employs a conventional approach that utilizes the principles of gradient descent through backpropagation. The goal is to adjust the node features in a manner that directly contributes to the two adversarial objectives. Each iteration of feature optimization is guided by the following rule:
$F_{s}=F_{s}-lr\cdot\nabla_{F_{s}}\mathcal{L}_{atk}$ (10)
4.2.2 Structure Optimization
$\nabla_{str}=\nabla_{A_{p}}\mathcal{L}_{atk}^{link}+\beta\cdot\nabla_{\hat{A}}\mathcal{L}_{atk}^{cls}$ (11)
In this formula, $\nabla_{A_{p}}\mathcal{L}_{atk}^{link}$ denotes the gradient of attack loss by the link prediction model with respect to the subgraph adjacency matrix $A_{p}$ before link recommender intervention, while $\nabla_{\hat{A}}\mathcal{L}_{atk}^{cls}$ represents the gradient of attack loss by the node classification model with respect to the post-update adjacency matrix $\hat{A}$ modified by the link recommender. The parameter $\beta$ acts as a balancing factor in the optimization of the structure. The structural update process involves ranking the combined gradient $\nabla_{str}$, then strategically adding and removing edges based on this ranking. Specifically, we add edges in locations where they are currently absent, prioritizing those with the lowest combined gradient values. Simultaneously, we remove edges from areas with the highest combined gradient values.
5 Experiments
5.1 Experiment Settings
5.1.1 Datasets
To evaluate the effectiveness of our attack methods, we perform experiments on five distinct datasets. These datasets include Cora and PubMed, both citation networks, Amazon-Photo and Computers, co-purchase networks from a popular e-commerce platform, and FacebookPagePage, representing a social network graph.
5.1.2 Experimental Verification
We conduct an experimental verification of our LiSA framework using GAE and VGAE [8] for link recommender, along with GCN [9], SGC [15], and GraphSAGE [7] as node classifiers. Our attack methodology is applied across all datasets with the presumption that the target user would establish links with nodes of top three linking scores in the entire poisoned graph. To assess the feasibility of our methods, we randomly select 1000 nodes from each original graph as target victim nodes. We then generate subgraphs and execute attacks on these nodes individually. This approach allows us to compute both the link success rate (LSR) and the overall attack success rate (ASR). The LSR is defined as the proportion where at least one adversarial edge is generated between the target node and subgraph, while the ASR refers to the misclassfication rate on these victim nodes. For the link recommender, we use 85% of the edges for training, with half serving as positive edge labels with negative sampling for both subgraph optimization and attack verification. For node classification, 5% of the nodes are randomly chosen for training labels. These labels are utilized in subgraph optimization and the pretraining of a node classifier to identify the second highest output class as the attack label. To maintain a consistent basis for comparison across different datasets, we employ a subgraph comprising five nodes for each target victim node attack. However, we tailor hyperparameters such as the number of subgraph edges $n_{E}$, and the balance parameters $\alpha$ and $\beta$ for each dataset, ensuring optimal performance and adaptability to the specific characteristics of each dataset. We also compare the LiSA framework with baseline of GraphCopy in the SIA scenario, and an additional baseline of conventional NIA scenario. This particular version of NIA incorporates the probability of successful link creation, aiming to demonstrate the influence of link generation on the success of NIAs.
Figure 3: Attack outcome on GCN with SIA and NIA scenario. For SIA, the link recommender model is GAE.
5.2 Attack Results
The attack results using the LiSA framework are presented in Table 1, while the misclassification rates for the original clean graphs are also provided, serving as a reference for evaluating the impact of our attacks. The outcomes of the attack indicate that the subgraphs produced by the LiSA framework successfully establish links between target nodes and subgraphs, leading to a notable decrease in the accuracy of node classification performance. This demonstrates the effectiveness of the LiSA framework within the SIA scenario. As illustrated in Fig. 3, the LiSA framework surpasses GraphCopy by achieving a higher ASR, thereby demonstrating its superiority over GraphCopy as an optimization-based method. Notably, the average attack budget for GraphCopy is proportional to the square of the average node degree in the dataset. In denser graphs, such as Amazon-Photo and Computers, attacking a single node requires a subgraph size exceeding 10% of the original graph. Conversely, LiSA achieves higher ASR with a much smaller budget, requiring only five nodes for each target node attack, highlighting the efficiency and practicality in real-world scenarios.
Table 1: Attack results (LSR % | ASR %) and clean graph misclassification (%).
  Cora Photo Computers PubMed PagePage
  GCN 18.7 7.7 14.7 21.6 11.7
SGC 18.7 50.1 44.0 25.3 26.8
GraphSAGE 21.4 7.4 13.8 22.5 15.1
  GAE & GCN 52.3 | 48.6 31.0 | 30.4 34.0 | 35.5 82.3 | 68.5 50.4 | 48.0
GAE & SGC 52.1 | 56.1 34.7 | 55.5 37.7 | 43.5 79.9 | 68.8 51.6 | 61.6
GAE & GraphSAGE 52.6 | 55.1 34.5 | 29.2 35.8 | 32.6 80.2 | 64.3 49.6 | 47.0
VGAE & GCN 44.6 | 50.3 25.2 | 29.5 32.9 | 37.2 77.3 | 66.8 42.4 | 45.0
VGAE & SGC 45.9 | 59.9 26.2 | 58.2 33.9 | 43.1 79.5 | 68.2 35.8 | 52.6
VGAE & GraphSAGE 46.8 | 55.5 26.8 | 21.8 35.6 | 32.5 79.4 | 67.4 40.2 | 37.4
5.3 Impacts of Subgraph Size and User Added Links
We conduct a series of experiments on the Amazon-Photo dataset to investigate how subgraph size impact the attack. Newly generated links are chosen based on top linking scores in the experiments. Our initial observation reveal a decline in LSR as the subgraph size increased, a trend depicted in Fig. 4(a). This decline likely stems from the challenges associated with optimizing larger subgraphs, especially when adjusting both structure and features. However, we noted a stabilization in LSR when the subgraph size exceeds 5. Another interesting finding related to the response of ASR to the number of links added by users, as illustrated in Fig. 4(b). With the addition of a single link, we observe that the ASR is highest when the subgraph consist of a single node, decreasing as the subgraph size expands, possibly due to the corresponding reduction in LSR. On the other hand, with five links added, the ASR improve as the subgraph grow to five nodes, before experiencing a slight decrease with further increases in subgraph size. This phenomenon can be attributed to the rise in aggressor nodes linked to the target, enhancing effectiveness of the attack. However, exceeding the subgraph size with available links prompts the link recommender to seek top-scoring nodes from the original graph for link formation, which strengthens the resilience of target node to the attack [3]. Taking into account these effects, we found that the optimal ASR of 10 links was attained when the subgraph size was 7.
(a) LSR (%)
(b) ASR (%)
Figure 4: Impacts of subgraph size and user added links on Amazon-Photo.
5.4 Subgraph Analysis
(a) Cora
(b) PubMed
Figure 5: Subgraph Analysis.
5.5 Ablation Studies
We perform ablation studies of LiSA on the Cora and PubMed dataset as depicted in Fig. 6. This analysis involves experiments with variations of the LiSA framework, including the omission of the surrogate models for node classification and link prediction (denoted as w/o cls and w/o link, respectively). Additionally, we explore the impact of focusing solely on node features optimization (labeled as w/o str) and solely on adjacency matrix optimization with randomly initialized node features (labeled as w/o feat) to discern the specific contributions of these optimization strategies.
(a) Cora
(b) PubMed
Figure 6: Ablation studies.
These findings reveal that adversarial links can be formed effectively through the use of the link prediction surrogate model, with the subgraph thus generated showing a marginal ASR improvement against the clean graph. Conversely, relying only on the node classification surrogate model fails to generate any links between target node and the subgraph. This highlights the significance of utilizing dual surrogate models, which can greatly enhance the attack to accomplish both adversarial objectives. Regarding optimization strategies, we observe that focusing exclusively on the structural aspects without adjusting the node features has no improvement on the attack. However, optimizing only the node features partially achieve the adversarial objectives, indicating that feature optimization plays a primary role in the SIA scenario. Nevertheless, structural optimization emerges as a critical supporting factor, enhancing the overall efficacy of the attack. Thus, the experimental results suggest that each component of LiSA is crucial to improve the attack performance.
6 Conclusions
This study presents a new attack scenario, along with a novel attack framework that utilizes link recommender to compromise the node classification accuracy of GNNs through subgraph injection. Unlike conventional attacks that require direct manipulation of the original graph or the addition of links to artificially created nodes, LiSA executes an attack indirectly by injecting an isolated subgraph, thereby enhancing practicality in real-world scenarios. Our experiments across various real-world datasets demonstrate the effectiveness of LiSA, illustrating significant degradation in GNN performance.
{credits}
6.0.1 Acknowledgements
Appendix 0.A Training Algorithm
The training algorithm of LiSA is shown in Algorithm 1.
@@ Appendix 0.A Training Algorithm
Appendix 0.B GraphCopy Settings and Attack Results
GraphCopy introduces a fundamental approach for conducting adversarial attacks, achieving the adversarial objectives to a certain degree, as demonstrated in Table 2. However, it faces two significant limitations that reduce its effectiveness and practicality in real-world scenarios. First, simply duplicating the graph and altering node features does not ensure a high success rate for adversarial attacks. Second, replicating the n-hop neighborhood of a target node leads to a subgraph size that increases quadratically with the node degree. For nodes with high degrees, the attack budget increase substantially, making GraphCopy impractical for use in dense or large-scale networks. In the experimental configuration of GraphCopy, we create the subgraph by replicating the 2-hop neighborhoods surrounding the target nodes and introducing a 10% Gaussian perturbation to the features of the duplicated nodes.
@@ Appendix 0.B GraphCopy Settings and Attack Results
Table 2: GraphCopy attack results (LSR % | ASR %) and clean graph misclassification rate (%).
  Cora Photo Computers PubMed PagePage
  GCN 18.7 7.7 14.7 21.6 11.7
@@
  GAE & GCN 13.3 | 25.8 30.5 | 17.7 45.3 | 29.1 61.8 | 44.2 27.4 | 30.8
GAE & SGC 14.0 | 35.5 30.8 | 56.0 44.4 | 46.3 62.6 | 51.7 29.0 | 56.0
GAE & GraphSAGE 15.3 | 33.0 31.5 | 12.9 43.9 | 17.8 63.3 | 41.1 27.6 | 20.6
VGAE & GCN 29.9 | 36.8 40.7 | 17.6 48.8 | 29.8 64.1 | 44.7 50.0 | 25.4
VGAE & SGC 30.3 | 42.6 39.4 | 56.8 50.9 | 45.9 65.1 | 49.7 44.8 | 50.2
VGAE & GraphSAGE 31.1 | 39.9 39.8 | 14.1 50.8 | 20.5 63.6 | 42.7 43.6 | 23.0
Appendix 0.C Dataset Statistics
@@
Table 3: Dataset statistics.
@@
Photo 7,650 238,162 31.13 746 8
Computers 13,752 491,722 35.76 767 10
@@
PagePage 22,470 342,004 15.22 128 4
Appendix 0.D NIA Settings, Results and Analysis